A current of 10A is flowing in a wire of lenght 1.5m.A force of 15N acts on it when it is placed in a uniform magnetic field of 2T. What is the angle between magnetic field and  the direction of current??
 

Grade:12

1 Answers

HARI CHANDANA
6 years ago
Given,
i=10 A,
length=1.5m,
force=15N,
magnetic field [B]=2T,
we need to find angle..
  • we have,F=Bilsin​\theta​​
  • substituting 15=2.×10×1.5sin\Theta
  • sin\Theta=1/2,
  • \Theta=30 degrees.